# The Illusion of Consistency: Selection-Induced Bias in Gated Kalman Innovation Statistics

### v1

@@ Abstract
This paper shows that innovation statistics computed after gating converge to gate-conditioned rather than nominal quantities. Under classical linear–Gaussian assumptions, we derive exact expressions for the first- and second-order moments of the innovation conditioned on ellipsoidal gating, and show that gating induces a deterministic, dimension-dependent contraction of the innovation covariance.
The analysis is extended to NN association, which is shown to act as an additional statistical selection operator. We prove that selecting the minimum-norm innovation among multiple in-gate measurements introduces an unavoidable energy contraction, implying that nominal innovation statistics cannot be preserved under nontrivial gating and association. Closed-form results in the two-dimensional case quantify the combined effects and illustrate their practical significance.
@@ I Introduction
The objective of this paper is to isolate and characterize the statistical effects induced solely by validation gating and NN association, independent of clutter, false alarms, nonlinear dynamics, or modeling errors. Within the classical linear–Gaussian Kalman filtering framework, we provide an exact characterization of the innovation moments conditioned on ellipsoidal gating and show that validation gating induces a deterministic, dimension-dependent contraction of the innovation covariance. We further show that NN association acts as an additional statistical selection operator and introduces an unavoidable, multiplicity-dependent contraction of innovation energy through order-statistic selection. Together, these results imply that nominal innovation statistics cannot be preserved under nontrivial gating and association, even when all Kalman filter assumptions are satisfied.
@@ I-A Related Work
Innovation-based statistics are a fundamental component of Kalman filtering theory and practice, where quantities such as the NIS are routinely used for consistency monitoring, measurement validation, and adaptive tuning. Under nominal linear Gaussian assumptions, the statistical properties of the innovation process are well understood and documented in classical estimation references [kalman1960new, kalman1961new, jazwinski2007stochastic, bar2004estimation, brown1992introduction].
In practical tracking systems, innovation statistics are commonly employed in adaptive filtering and noise covariance estimation schemes, often relying on assumed chi-square properties of the NIS [mehra1970identification, akhlaghi2017adaptive, wang2018variational, youn2020novel]. Such approaches implicitly assume that the observed innovation sequence is representative of the nominal innovation distribution. However, in operational settings, innovation statistics are almost always evaluated after validation gating and data association, conditions under which these assumptions may no longer hold.
Validation gating and NN association are standard components of tracking systems in aerospace, radar, and navigation applications [bar2004estimation, ramachandra2018kalman, shats1988discrete, karlsson2017future]. Despite their widespread use, the statistical impact of these selection mechanisms on innovation-based diagnostics has received comparatively little explicit theoretical treatment.
@@
Recent work on adaptive and learning enhanced Kalman filtering further highlights the reliance on innovation statistics for online consistency assessment and tuning [or2021kalman, or2022hybrid, or2023learning]. These methods benefit from a precise understanding of how structural elements of the tracking pipeline, such as gating and association, affect the observed innovation process.
@@ III-C Gate-Conditioned NIS Statistics
Figure 1: Gate-conditioned NIS distribution in the two-dimensional case. The nominal $\chi^{2}_{2}$ distribution is shown together with the truncated distribution induced by ellipsoidal validation gating. The gate-conditioned mean is systematically lower than the nominal reference value, illustrating the apparent overconfidence induced by gating even under ideal Kalman filter assumptions.
@@ III-D Implications for Innovation-Based Diagnostics
Any diagnostic or adaptive tuning procedure operating on post-gate innovation statistics is therefore estimating gate-conditioned quantities. Interpreting such statistics as nominal systematically leads to apparent overconfidence, even in correctly tuned Kalman filters.
@@ IV-A Classical use of NIS for consistency Validation
Under the nominal Kalman filter assumptions recalled in Section III, the NIS follows a chi-square distribution with $m$ degrees of freedom. As a result, standard practice is to compare instantaneous NIS values to chi-square confidence bounds, or to monitor the empirical mean of $\{Z_{k}\}$ over time and compare it to the nominal reference value $m$, as commonly done in online consistency testing [piche2016online].
@@ IV-B Post-gate NIS is not chi-square
but instead follows a truncated chi-square law obtained by renormalizing $\chi^{2}_{m}$ on the interval $0,\tau] [$dunik2020covariance].
@@ IV-C Gate-aware interpretation of NIS statistics
This normalization allows classical chi-square intuition to be reused without modifying the underlying Kalman filter or gating logic, while preserving consistency interpretations used in practice [piche2016online].
@@ IV-D Implications for adaptive tuning and diagnostics
Many adaptive noise-tuning and consistency-monitoring schemes rely, either explicitly or implicitly, on innovation covariance estimates or NIS-based statistics [dunik2017noise], and often enforce nominal chi-square innovation behavior as a tuning objective [chen2023kalman].
@@ V Bias Induced by Nearest-Neighbor Association
The previous sections characterized the statistical effect of ellipsoidal gating on innovation moments. In practical tracking systems, however, gating is only the first stage of a selection process. When multiple measurements fall inside the validation region, a data association rule is applied to select a single measurement for the state update. The most widely used rule in classical tracking is NN association [reid2003algorithm, bar1988tracking].
@@ V-A Nearest-neighbor association as statistical selection
Consider a time step at which $M\geq 1$ measurements pass the validation gate $\mathcal{A}$. Let $\{\nu^{(i)}\}_{i=1}^{M}$ denote the corresponding innovation vectors, each generated according to the same post-gate distribution $(\nu\mid\mathcal{A})$. NN association selects the innovation with minimum NIS, as originally proposed in the classical multi-target tracking framework of Reid [reid2003algorithm],
@@ V-E Implications for innovation-based diagnostics
The bound (18) has direct implications for innovation-based diagnostics and tuning. Statistics computed from NN-selected innovations exhibit systematically reduced energy relative to both nominal and gate-conditioned references. Consequently, diagnostic rules based on expected NIS values will indicate apparent overconfidence even when the system operates exactly according to the nominal model.
Crucially, this behavior does not arise from clutter, false alarms, or modeling errors. It is a structural consequence of NN association itself. As such, it cannot be eliminated by tuning and must be explicitly accounted for when interpreting innovation statistics in gated tracking systems.
@@ VI-D Interpretation
The two-dimensional case provides a concrete numerical instantiation of the general results derived earlier. In this setting, the effect of ellipsoidal gating can be quantified explicitly, revealing a predictable contraction of innovation energy that depends only on the gate probability. NN association further compounds this effect through order statistic selection. As illustrated in Table I, even commonly used gate probabilities lead to a substantial reduction in post-gate innovation energy. Consequently, innovation-based diagnostics computed after gating or NN association, when interpreted using nominal reference values, systematically indicate apparent overconfidence.
@@ VII Discussion
The analysis developed in this paper provides a principled explanation for a widely observed phenomenon in Kalman-based tracking systems: the tendency of well-functioning trackers to exhibit apparent overconfidence when innovation-based diagnostics are evaluated after gating and association.
@@ VII-B Implications for implementation and evaluation
  • Apparent overconfidence in post-gate diagnostics does not necessarily indicate filter mis-tuning.
@@ A-D Bounds on $\gamma(\tau,m)$
[Uncaptioned image] Barak Or (Member, IEEE) received a B.Sc. degree in aerospace engineering (2016), a B.A. degree (cum laude) in economics and management (2016), and an M.Sc. degree in aerospace engineering (2018) from the Technion–Israel Institute of Technology. He graduated with a Ph.D. degree from the University of Haifa, Haifa (2022). His research interests include navigation, deep learning, sensor fusion, and estimation theory.

### v2

# Selection-Induced Contraction of Innovation Statistics in Gated Kalman Filters

@@ Abstract
This paper shows that innovation statistics computed after gating converge to gate-conditioned rather than unconditional nominal reference quantities. Under classical linear-Gaussian assumptions, we derive exact expressions for the first- and second-order moments of the innovation conditioned on ellipsoidal gating, and show that gating induces a deterministic, dimension-dependent contraction of the innovation covariance relative to the nominal reference.
The analysis is extended to nearest-neighbor (NN) association, which is shown to act as an additional statistical selection operator. We prove that selecting the minimum norm innovation among multiple in-gate measurements introduces an unavoidable energy contraction, implying that nominal innovation reference statistics cannot be preserved under nontrivial gating and association due to deterministic selection effects, even under perfectly matched linear-Gaussian Kalman filter assumptions. Closed-form results in the two-dimensional case quantify the combined effects and illustrate their practical significance.
@@ I Introduction
The objective of this paper is to isolate and characterize the statistical effects induced solely by validation gating and NN association, independent of clutter, false alarms, nonlinear dynamics, or modeling errors. Within the classical linear–Gaussian Kalman filtering framework, we provide an exact characterization of the innovation moments conditioned on ellipsoidal gating and show that validation gating induces a deterministic, dimension-dependent contraction of the innovation covariance relative to the nominal reference. We further show that NN association acts as an additional statistical selection operator and introduces an unavoidable, multiplicity-dependent contraction of innovation energy through order-statistic selection. Together, these results imply that nominal innovation statistics cannot be preserved under nontrivial gating and association, even when all Kalman filter assumptions are satisfied.
@@ I-A Related Work
Innovation-based statistics are a fundamental component of Kalman filtering theory and practice, where quantities such as the NIS are routinely used for consistency monitoring, measurement validation, and adaptive tuning. Under nominal linear Gaussian assumptions, the statistical properties of the innovation process are well understood and documented in classical estimation references [10, 9, 8, 3, 4].
In practical tracking systems, innovation statistics are commonly employed in adaptive filtering and noise covariance estimation schemes, often relying on assumed chi-square properties of the NIS [12, 1, 20, 21]. Such approaches implicitly assume that the observed innovation sequence is representative of the nominal innovation distribution. However, in operational settings, innovation statistics are almost always evaluated after validation gating and data association, conditions under which these assumptions may no longer hold.
Validation gating and NN association are standard components of tracking systems in aerospace, radar, and navigation applications [3, 17, 19, 11]. Despite their widespread use, the statistical impact of these selection mechanisms on innovation-based diagnostics has received comparatively little explicit theoretical treatment.
@@
Recent work on adaptive and learning enhanced Kalman filtering further highlights the reliance on innovation statistics for online consistency assessment and tuning [13, 14, 15]. These methods benefit from a precise understanding of how structural elements of the tracking pipeline, such as gating and association, affect the observed innovation process.
@@ III-C Gate-Conditioned NIS Statistics
Figure 1: Gate-conditioned NIS distribution in the two-dimensional case. The nominal $\chi^{2}_{2}$ distribution is shown together with the truncated distribution induced by ellipsoidal validation gating. The gate-conditioned mean is systematically lower than the nominal reference value, illustrating the systematic contraction of innovation energy relative to the nominal reference induced by validation gating, even under ideal Kalman filter assumptions.
@@ III-D Implications for Innovation-Based Diagnostics
Any diagnostic or adaptive tuning procedure operating on post-gate innovation statistics is therefore estimating gate-conditioned quantities. Interpreting post-selection innovation statistics using nominal, unconditional reference values systematically leads to the appearance of reduced innovation energy relative to the nominal covariance.
When nearest-neighbor association is applied, this effect is further amplified by order-statistic selection and becomes pronounced in high-density scenarios, precisely in regimes where NN association is known to be theoretically suboptimal.
@@ IV-A Classical use of NIS for consistency Validation
Under the nominal Kalman filter assumptions recalled in Section III, the NIS follows a chi-square distribution with $m$ degrees of freedom. As a result, standard practice is to compare instantaneous NIS values to chi-square confidence bounds, or to monitor the empirical mean of $\{Z_{k}\}$ over time and compare it to the nominal reference value $m$, as commonly done in online consistency testing [16].
@@ IV-B Post-gate NIS is not chi-square
but instead follows a truncated chi-square law obtained by renormalizing $\chi^{2}_{m}$ on the interval $0,\tau] [$6].
@@ IV-C Gate-aware interpretation of NIS statistics
This normalization allows classical chi-square intuition to be reused without modifying the underlying Kalman filter or gating logic, while preserving consistency interpretations used in practice [16].
@@ IV-D Implications for adaptive tuning and diagnostics
Many adaptive noise-tuning and consistency-monitoring schemes rely, either explicitly or implicitly, on innovation covariance estimates or NIS-based statistics [7], and often enforce nominal chi-square innovation behavior as a tuning objective [5].
@@ V Bias Induced by Nearest-Neighbor Association
The previous sections characterized the statistical effect of ellipsoidal gating on innovation moments. In practical tracking systems, however, gating is only the first stage of a selection process. When multiple measurements fall inside the validation region, a data association rule is applied to select a single measurement for the state update. The most widely used rule in classical tracking is NN association [18, 2].
@@ V-A Nearest-neighbor association as statistical selection
Consider a time step at which $M\geq 1$ measurements pass the validation gate $\mathcal{A}$. Let $\{\nu^{(i)}\}_{i=1}^{M}$ denote the corresponding innovation vectors, each generated according to the same post-gate distribution $(\nu\mid\mathcal{A})$. NN association selects the innovation with minimum NIS, as originally proposed in the classical multi-target tracking framework of Reid [18],
@@ V-E Implications for innovation-based diagnostics
The bound (18) has direct implications for innovation-based diagnostics and tuning. Statistics computed from NN-selected innovations exhibit systematically reduced energy relative to both nominal and gate-conditioned references. Consequently, diagnostic rules based on expected NIS values will indicate a systematic reduction of innovation energy relative to nominal reference values, even when the system operates exactly according to the nominal model. Crucially, this behavior does not arise from clutter, false alarms, or modeling errors. It is a structural consequence of NN association itself. As such, it cannot be eliminated by tuning and must be explicitly accounted for when interpreting innovation statistics in gated tracking systems.
@@ VI-D Interpretation
The two-dimensional case provides a concrete numerical instantiation of the general results derived earlier. In this setting, the effect of ellipsoidal gating can be quantified explicitly, revealing a predictable contraction of innovation energy that depends only on the gate probability. NN association further compounds this effect through order statistic selection. As illustrated in Table I, even commonly used gate probabilities lead to a substantial reduction in post-gate innovation energy. Consequently, innovation-based diagnostics computed after gating or NN association, when interpreted using nominal reference values, systematically indicate a reduction of innovation energy relative to the nominal reference.
@@ VII Discussion
The analysis developed in this paper provides a principled explanation for a widely observed phenomenon in Kalman-based tracking systems: the tendency of innovation-based diagnostics evaluated after gating and association to exhibit systematically reduced innovation energy relative to nominal reference values, even in well-functioning trackers.
@@ VII-B Implications for implementation and evaluation
  • Systematically reduced innovation energy observed in post-gate diagnostics does not necessarily indicate filter mis-tuning.
@@ VIII Conclusion
Acknowledgment
The author gratefully acknowledges Prof. Yaakov Bar-Shalom for his insightful comments that helped clarify the interpretation and terminology of the paper.